OpenStudy (anonymous):

need help using scientific and engineering notation. these are the problems i'm working on: 1.2 x 10e4 = ? x 10e3 0.5 x 10e-4 = ? x 10e-6 125.0 x 10e5 = ? x 10e6 250 x 10e-5 = ? x 10e-4

OpenStudy (amistre64):

"e" in this light i believe means "exponent"?

OpenStudy (anonymous):

yeah i'm dealing with powers of ten

OpenStudy (amistre64):

1234 (10^2) = 123400 , which tells me the decimal moves right when we increase the exponent. and by default, moves left when it decreases

OpenStudy (anonymous):

\[10^{4}=1000\]\[1.2\times10^{4}=1200\]\[10^{3}=100\]\[1200=100\times?\]It really helps to visualize what they are in expanded form.

OpenStudy (amistre64):

1.2 x 10e4 = ? x 10e3 ; the exponent decreased by 1, so lets move the deci left by 1 1.2 x 10e4 = .12 x 10e3 id say

OpenStudy (anonymous):

Bah 1200 should be 12000

OpenStudy (amistre64):

0.5 x 10e-4 = ? x 10e-6 is moved 2 to the left 0.5 x 10e-4 = .005 x 10e-6 perhaps?

OpenStudy (anonymous):

ok so 1.2 times 10^4 would be 0.12 times 10^3?

OpenStudy (amistre64):

looks good to me

OpenStudy (amistre64):

... but id have to dbl chk myself to make sure ....

OpenStudy (anonymous):

so to convert them i just have to figure out where to move the decimal point?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

and the movement of the exponent is the clue

OpenStudy (amistre64):

1.00 * 10^0 = 1.00 1.00 * 10^1 = 10.0 1.00 * 10^2 = 100.

OpenStudy (anonymous):

so 0.5 times 10^-4 would be .005 times 10^-6?

OpenStudy (amistre64):

im sure it is, but let me do one more check to clear the cobwebs .. 1.00 * 10^1 = ? * 10^2 10 = ? * 100 10 = .1 * 100 ; sooo..... 1.00 * 10^1 = .1 * 10^2 yep, i had it backwards ...

OpenStudy (amistre64):

0.5 times 10^-4 = 50. times 10^-6 then

OpenStudy (anonymous):

ok thanks.

OpenStudy (amistre64):

when in doubt, just work it out with what you can remember till you recall what needs to be done :) good luck with it